OpenStudy (anonymous):

Find an equation of the parabola which passes through the points (1,3) and (-2,0). Please do it step-by-step...

OpenStudy (anonymous):

Basic form is y^2 = 4ax. This parabola goes through x is -2 so basic form is y^2 = 4a(x+2) Now use (1,3) to get a 9 = 4a(3) -> a=3/4 so y^2 = 3(x+2)

OpenStudy (amistre64):

you need 3 points to fix a parabola to a plane

OpenStudy (amistre64):

and we prolly need to assume its not a rotatation

OpenStudy (anonymous):

In general that's true but what else can one do given the question, except assume a basic parabola in standard position and then translate it.

OpenStudy (anonymous):

there are infinite parabolas passing through these points

OpenStudy (amistre64):

oh i dont deny we need to make some sort of assumption; but the question itself is messed up. There is no "correct" answer, even tho there are plenty of wrong answers :)